A 2.75 x 10–3 M Ba(OH)2 solution is prepared.

a. What is the pOH of the solution? Blank 1 ( write answer to the hundredths place)

b. What is the pH of the solution? Blank 2 ( write answer to the hundredths place)

c. Is the solution acidic, basic or neutral? Blank 3
Answer #1

Given:

[Ba(OH)2] = 2.75*10^-3 M

So,

[OH-] = 2*[Ba(OH)2]

= 2*2.75*10^-3 M

= 5.50*10^-3 M

a)

use:

pOH = -log [OH-]

= -log (5.5*10^-3)

= 2.26

Answer: 2.26

b)

use:

PH = 14 - pOH

= 14 - 2.26

= 11.74

Answer: 11.74

c)

Since pOH > pH, this is basic

Answer: Basic